# Phonon effects, impact ionization and power conversion in Mott photovoltaic systems

### v1

@@ Abstract
We analyze the effect of acoustic phonons on the photocurrent and the spectral characteristics of a simplified photovoltaic setup made of Mott insulating layers between two metallic leads among which a bias voltage is applied. This paper extends and completes the study made in a previous work by the authors [Gazzaneo et al., Phys. Rev. B 109, 235134]. We include acoustic phonons via the Migdal approximation and we use real-space Floquet dynamical mean-field theory to address the nonequilibrium Floquet steady-state. The so-called auxiliary master equation approach is employed as impurity solver. We find that impact ionization is only weakly affected by dissipation by acoustic phonons at low bias voltages. For higher biases instead, the Hartree shift considerably alters the on-site energies of the Hubbard bands and suppresses the photocurrent for intermediate electron-phonon coupling strengths. Impact ionization processes play a fundamental role in enhancing the electrical output power, which decreases when electron-phonon interaction is considered.
I Introduction
@@
At the nonequilibrium steady-state (NESS), the operating regime of a solar cell, heterostructures made of multiple correlated layers can be theoretically investigated via real-space Floquet dynamical mean-field theory (DMFT) [33, 34, 35, 36, 37, 22, 38, 39, 40, 41, 42]. In this regard, a tentative evaluation of the efficiency, which has been estimated by Manousakis [11] and evaluated by Petocchi [24] for times up to $\sim 100$ fs, would provide an important assessment for future investigations in this field. Moreover, despite the predominance of electron-electron interaction, electron-phonon (e-ph) scattering is an important dissipation mechanism in such systems, which competes with II as energy dissipation process of a high-energy photoexcited electron [14]. E-ph interaction modifies the density of states of the different layers, the overlap of which plays a fundamental role for transport properties, as shown in our previous work [31], with sizeable consequences on the photocurrent. In the present manuscript we therefore complement and generalize some results of Refs. [30, 31], including acoustic phonons  [^1] Acoustic phonons are more effective than optical phonons in energy dissipation. For this reason we concentrate our analysis on this case. via the non self-consistent Migdal approximation and analyzing the photocurrent-bias voltage characteristic at the NESS for a Mott photovoltaic device.
@@
Figure 1: The system considered in the manuscript. A central Mott insulating region which consists of $\text{L}=4$ correlated layers, in green, is irradiated with a periodic, monochromatic, light with frequency $\Omega$. In this region acoustic phonons are also included and contribute to dissipate the harvested energy from the external driving. Two uncorrelated interfaces act as contacts, in blue, and separate the correlated region from the metallic leads, in gray. A bias voltage $\Delta V$ applied to the leads keeps them at different chemical potentials, to collect the doublon and holons created by the external driving $\Omega$.
@@
The rest of the manuscript is organized as follows: in Sec. II we describe the model under investigation, in Sec. III we discuss the formalism and the mathematical tools hereby employed. We discuss the results in Sec. IV and in Sec. V we make our conclusions and final remarks.
II Model
We consider the system of Ref. [31] with the addition of acoustic phonons. The heterostructure is made up of L correlated layers arranged along the $z$-axis, i.e. $z\in\left\{1,...,\text{L}\right\}$, translationally invariant in the $xy$ plane, between two metallic leads ($\rho=\text{l},\text{r}$) as shown in Fig. 1.
@@ II Model
$\begin{split}\hat{H}(t)&=-\sum_{z,\langle{\bf r},{\bf r}^{\prime}\rangle,\sigma}t_{z}(t)\hat{c}_{z,{\bf r},\sigma}^{\dagger}\hat{c}_{z,{\bf r}^{\prime},\sigma}-\sum_{\langle z,z^{\prime}\rangle,{\bf r},\sigma}t_{zz^{\prime}}\hat{c}_{z,{\bf r},\sigma}^{\dagger}\hat{c}_{z^{\prime},{\bf r},\sigma}\\ &+\sum_{z,{\bf r}}U_{z}\hat{n}_{z,{\bf r},\uparrow}\hat{n}_{z,{\bf r},\downarrow}+\sum_{z,{\bf r},\sigma}\varepsilon_{z}^{(0)}\hat{n}_{z,{\bf r},\sigma}+\hat{H}_{\text{leads}}\\ &+\hat{H}_{\text{e-ph}}+\hat{H}_{\text{ph}}.\end{split}$ (1)
The operator $\hat{c}_{z,{\bf r},\sigma}^{\dagger}$ ($\hat{c}_{z,{\bf r},\sigma}$) is the creation(annihilation) operator of an electron on site ${\bf r}=(x,y)$ of layer $z$ of spin $\sigma=\{\uparrow,\downarrow\}$. $\hat{n}_{z,{\bf r},\sigma}=\hat{c}_{z,{\bf r},\sigma}^{\dagger}\hat{c}_{z,{\bf r},\sigma}^{\phantom{\dagger}}$ is the corresponding particle number operator. Here the brakets $\langle z,z^{\prime}\rangle$ indicate neighboring layers along the $z$-axis while $\langle{\bf r}^{\phantom{\dagger}},{\bf r}^{\prime}\rangle$ refers to neighboring sites belonging to the same layer.
The central correlated region is under a time-periodic, homogeneous and monochromatic, electric field with frequency $\Omega$, which enters in the Hamiltonian (1) via the Peierls substitution [46] in the intralayer hopping as
$t_{z}(t)=t_{z}\ e^{-\operatorname*{i}\frac{q}{\hbar}\left(\bm{r}-\bm{r}^{\prime}\right)\cdot\bm{A}(t)},$ (2)
where in Eq. (2) $\bm{A}$(t) is the time-dependent vector potential, $\hbar$ Planck’s constant and $q$ the charge of the electrons. We choose $\bm{A}(t)=A(t)\bm{e}_{0}$ to point along the lattice body diagonal of a hyper-cubic lattice $\bm{e}_{0}=(1,1,\dots,1)$, where $A(t)=\frac{\hbar}{qa}\mathscr{A}\sin(\Omega t)$ and $\mathscr{A}=-\frac{qE_{0}a}{\hbar\Omega}$, with $E_{0}$ is the electric field amplitude, and $a$ the lattice spacing [47, 48]. The electric field is then given in the temporal gauge by $\bm{E}=-\partial_{t}\bm{A}(t)=E_{0}\cos(\Omega t)\bm{e}_{0}$.
The second term in Eq. (1) describes the hopping of electrons in between layers with nearest-neighbor amplitude $t_{zz^{\prime}}$. The third term includes in the model the local onsite Hubbard interaction $U_{z}$ and the fourth one the onsite energies $\varepsilon_{z}^{(0)}$, the expression of which will be given in Sec. III.1. $\hat{H}_{\text{leads}}$ describes the Hamiltonian of the metallic leads, the details of which will be also given in Sec. III.1.
@@
$\hat{H}_{\text{e-ph}}=g\sum_{z,{\bf r},\sigma}\hat{n}_{z,{\bf r},\sigma}\hat{x}_{z,{\bf r}},$ (3)
where $g$ is the e-ph coupling constant and $\hat{x}_{z,{\bf r}}\equiv\frac{1}{\sqrt{2}}\left(\hat{b}^{\dagger}_{z,{\bf r}}+\hat{b}_{z,{\bf r}}\right)$, with $\hat{b}^{\dagger}_{z,{\bf r}}$ ($\hat{b}_{z,{\bf r}}$) which creates (annihilates) a phonon of the acoustic branch on site ${\bf r}=(x,y)$ of layer $z$, whose dispersion relation is described by $\hat{H}_{\text{ph}}$, discussed in Sec III.1.
@@
The xy-plane is modeled as a $d$-dimensional lattice in the limit $d\rightarrow\infty$ with the rescaling of the intralayer hopping as $t_{\parallel}=t_{\parallel}^{\ast}/(2\sqrt{d})$ [52]. Every momentum-dependent function depends on the electronic crystal momentum ${{\bf k}}$ only via $\epsilon=-2t_{\parallel}\sum_{i=1}^{d}\cos(k_{i}a)$ and $\overline{\epsilon}=-2t_{\parallel}\sum_{i=1}^{d}\sin(k_{i}a)$. Summations over the Brillouin zone are then performed using the joint density of states (JDOS) [47, 52] $\rho(\epsilon,\overline{\epsilon})=(1/\pi t_{\parallel}^{\ast 2})\exp[-(\epsilon^{2}+\overline{\epsilon}^{2})/t_{\parallel}^{\ast 2}]$.
In this paper we set $\hbar=k_{\text{B}}=a=1=q$ and $t_{\parallel}^{\ast}/2=1$ as unit of energy.
III Methods
III.1 Dyson equation
In this section we follow the derivation of Ref. [31].
@@ III.1 Dyson equation
$\begin{split}\underline{{\bf G}}^{-1}_{zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})&=\underline{{\bf G}}^{-1}_{0,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})-\underline{{\bf\Sigma}}_{zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})\\ &-\underline{{\bf\Sigma}}_{\textrm{e-ph},zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon}).\end{split}$ (4)
The quantities in Eq. (4) are matrices in the Floquet representation, which we use to describe the periodic NESS of the system [25, 30, 31]. In the rest of the paper, we denote a Floquet-represented matrix as either $X_{mn}$ or $\bf{X}$. The underline in Eq. (4) stands for the Keldysh structure
$\underline{{\bf X}}\equiv\begin{pmatrix}{\bf X}^{\text{R}}&{\bf X}^{\text{K}}\\ {\bf 0}&{\bf X}^{\text{A}}\\ \end{pmatrix},$ (5)
with $\textbf{X}^{\text{R,A,K}}$ being the retarded, advanced and Keldysh components. We recall that $\textbf{X}^{\text{A}}=(\textbf{X}^{\text{R}})^{\dagger}$ and $\textbf{X}^{\text{K}}\equiv\textbf{X}^{>}+\textbf{X}^{<}$, where $\textbf{X}^{\lessgtr}$ are the lesser and greater components [53, 54, 55, 56].
The GF which corresponds to the noninteracting terms in the Hamiltonian (1) reads
$\begin{split}&\underline{G}^{-1}_{0,mn,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})=\underline{g}^{-1}_{0,mn,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})\\ &-\left[v^{2}_{\text{l}}\underline{g}_{\text{l}}(\omega_{n};\epsilon)\delta_{z,1}+v^{2}_{\text{r}}\underline{g}_{\text{r}}(\omega_{n};\epsilon)\delta_{z,\text{L}}\right]\delta_{mn}\delta_{zz^{\prime}},\end{split}$ (6)
@@
$\begin{split}&\left[g^{-1}_{0}(\omega_{n};\epsilon,\overline{\epsilon}))\right]^{\text{R}}_{mn,zz^{\prime}}=\left(\omega_{n}+\operatorname*{i}0^{+}-\varepsilon_{z}\right)\delta_{mn}\delta_{zz^{\prime}}\\ &-t_{zz^{\prime}}\delta_{mn}-\varepsilon_{mn}(\epsilon,\overline{\epsilon})\delta_{zz^{\prime}},\end{split}$ (7)
@@
$\varepsilon_{z}=\varepsilon_{z}^{(0)}+\frac{\Phi}{2}-\frac{(z-1)\Phi}{(\text{L}-1)},$ (8)
with $\varepsilon_{z}^{(0)}=-U/2$ at half-filling. The off-diagonal Floquet terms in Eq. (LABEL:eq:inv_non-int_lat_GF_comps) are given by the Floquet dispersion relation $\varepsilon_{mn}$ for a periodic field in a hyper-cubic lattice [47]
$\varepsilon_{mn}(\epsilon,\overline{\epsilon})=\begin{cases}J_{m-n}(\mathscr{A})\ \epsilon&m-n:\text{even}\\ \operatorname*{i}J_{m-n}(\mathscr{A})\ \overline{\epsilon}&m-n:\text{odd},\end{cases}$ (9)
where $J_{n}$ is the $n$-th order Bessel function of the first kind with argument $\mathscr{A}$ defined as in Sec. II.
@@
$\begin{split}g^{\text{R}}_{\rho}(\omega;\epsilon)&=\left(\omega-\varepsilon_{\rho}(\epsilon)+\operatorname*{i}\gamma_{\rho}\right)^{-1},\\ g^{\text{K}}_{\rho}(\omega;\epsilon)&=[g_{\rho}^{\text{R}}(\omega;\epsilon)-g_{\rho}^{\text{A}}(\omega;\epsilon)][1-2f(\omega,\mu_{\rho},\beta)],\end{split}$ (10)
@@
The electron self-energy (SE) $\underline{\bf{\Sigma}}_{zz^{\prime}}$ and the e-ph SE $\underline{{\bf\Sigma}}_{\textrm{e-ph},zz^{\prime}}$ are obtained from DMFT [57, 58, 59] and its nonequibrium extension for driven inhomogeneous systems, i.e. real-space Floquet DMFT (F-DMFT) [47, 60, 33, 22, 36, 41, 42]. In this approximation, they are independent of the crystal momentum and spatially local, i.e. $\underline{{\bf\Sigma}}_{zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{\bf{\Sigma}}_{z}(\omega)\delta_{zz^{\prime}}$, $\underline{{\bf\Sigma}}_{\textrm{e-ph},zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{{\bf\Sigma}}_{\textrm{e-ph},z}(\omega)\delta_{zz^{\prime}}$. Details regarding real-space F-DMFT are given in Appendix A.
@@
$\Sigma_{\text{e-ph},z}(z,z^{\prime})=\operatorname*{i}g^{2}G_{\text{loc},zz}(z,z^{\prime})D_{\text{ph}}(z,z^{\prime}).$ (11)
One can easily extract the retarded and Keldysh components of the e-ph SE from Eq. (11), which are listed in Ref. [49].
In the current setup, away from half-filling, the Retarded component of $\underline{\bm{\Sigma}}_{\text{e-ph},z}(\omega)$ has also the contribution of the so-called Hartree shift $\varepsilon_{\textrm{H},z}$, which enters in the $(0,0)$-Floquet matrix element as
$\Sigma_{\text{e-ph},z,00}^{\text{R}}(\omega)=\varepsilon_{\textrm{H},z}+\Sigma_{\text{e-ph},z,00,\textrm{XC}}^{\text{R}}(\omega),$ (12)
where $\Sigma_{\text{e-ph},z,00,\textrm{XC}}^{\text{R}}(\omega)$ is the contribution coming from Eq. (11). The expression of $\varepsilon_{\textrm{H},z}$ and its derivation are given in Sec. IV.1 and Appendix B respectively.
The Retarded component of the non-interacting phonon GF $\underline{D}_{\text{ph}}(t,t^{\prime})$ is given by [63]
$D^{\text{R}}_{\text{ph}}(t,t^{\prime})=-\operatorname*{i}\theta(t-t^{\prime})\int\mathop{}\!\mathrm{d}\omega\ \mathrm{e}^{-\operatorname*{i}\omega\left(t-t^{\prime}\right)}A_{\text{ph}}(\omega),$ (13)
where $A_{\text{ph}}(\omega)=(\omega/\omega^{2}_{\text{ph}})\mathrm{e}^{-|\omega|/\omega_{\text{ph}}}$ is the acoustic phonons’ spectral function, with $\omega_{\text{ph}}$ a soft cutoff frequency [64, 30]. The other Keldysh components are reconstructed via the fluctuation-dissipation theorem [65, 30].
III.2 Observables

### v2

@@ Abstract
We analyze the effect of acoustic phonons on the photocurrent and the spectral characteristics of a simplified photovoltaic setup made of Mott insulating layers between two metallic leads among which a bias voltage is applied. We include acoustic phonons via the Migdal approximation and we use real-space Floquet dynamical mean-field theory to address the nonequilibrium Floquet steady-state. The so-called auxiliary master equation approach is employed as impurity solver. We find that impact ionization is only weakly affected by dissipation by acoustic phonons at low bias voltages. For higher biases instead, the Hartree shift considerably alters the on-site energies of the Hubbard bands and suppresses the photocurrent for intermediate electron-phonon coupling strengths. Impact ionization processes play a fundamental role in enhancing the electrical output power, which decreases when electron-phonon interaction is considered.
1 Introduction
@@
At the nonequilibrium steady-state (NESS), the operating regime of a solar cell, heterostructures made of multiple correlated layers can be theoretically investigated via real-space Floquet dynamical mean-field theory (DMFT) [33, 34, 35, 36, 37, 22, 38, 39, 40, 41, 42]. In this regard, a tentative evaluation of the efficiency, which has been estimated by Manousakis [11] and evaluated by Petocchi [24] for times up to $\sim 100$ fs, would provide an important assessment for future investigations in this field. Moreover, despite the predominance of electron-electron interaction, electron-phonon (e-ph) scattering is an important dissipation mechanism in such systems, which competes with II as energy dissipation process of a high-energy photoexcited electron [14]. E-ph interaction modifies the density of states of the different layers, the overlap of which plays a fundamental role for transport properties [36, 37], with sizeable consequences on the photocurrent. In the present manuscript we therefore include acoustic phonons  [^1] Acoustic phonons are more effective than optical phonons in energy dissipation. For this reason we concentrate our analysis on this case. via the non self-consistent Migdal approximation and analyze the photocurrent-bias voltage characteristic at the NESS for a Mott photovoltaic device.
@@
Figure 1: The system considered in the manuscript. A central Mott insulating region which consists of $\mathrm{L}=4$ correlated layers, in green, is irradiated with a periodic, monochromatic, light with frequency $\Omega$. In this region acoustic phonons are also included and contribute to dissipate the harvested energy from the external driving. Two uncorrelated interfaces act as contacts, in blue, and separate the correlated region from the metallic leads, in gray. A bias voltage $\Delta V$ applied to the leads keeps them at different chemical potentials, to collect the doublon and holons created by the external driving $\Omega$.
@@
The rest of the manuscript is organized as follows: in Sec. 2 we describe the model under investigation, in Sec. 3 we discuss the formalism and the mathematical tools hereby employed. We discuss the results in Sec. 4 and in Sec. 5 we make our conclusions and final remarks.
2 Model
We consider the system shown in Fig. 1, given by a heterostructure made up of $\mathrm{L}$ correlated layers arranged along the $z$-axis, i.e. $z\in\left\{1,...,\mathrm{L}\right\}$, translationally invariant in the $xy$ plane, between two metallic leads ($\rho=\mathrm{l},\mathrm{r}$).
@@ 2 Model
$\hat{H}(t)$ $=-\sum_{z,\langle{\mathbf{r}},{\mathbf{r}}^{\prime}\rangle,\sigma}t_{z}(t)\hat{c}_{z,{\mathbf{r}},\sigma}^{\dagger}\hat{c}_{z,{\mathbf{r}}^{\prime},\sigma}-\sum_{\langle z,z^{\prime}\rangle,{\mathbf{r}},\sigma}t_{zz^{\prime}}\hat{c}_{z,{\mathbf{r}},\sigma}^{\dagger}\hat{c}_{z^{\prime},{\mathbf{r}},\sigma}$ (1)
$+\sum_{z,{\mathbf{r}}}U_{z}\hat{n}_{z,{\mathbf{r}},\uparrow}\hat{n}_{z,{\mathbf{r}},\downarrow}+\sum_{z,{\mathbf{r}},\sigma}\varepsilon_{z}^{(0)}\hat{n}_{z,{\mathbf{r}},\sigma}+\hat{H}_{\mathrm{leads}}$
$+\hat{H}_{\mathrm{e-ph}}+\hat{H}_{\mathrm{ph}}.$
The operator $\hat{c}_{z,{\mathbf{r}},\sigma}^{\dagger}$ ($\hat{c}_{z,{\mathbf{r}},\sigma}$) is the creation(annihilation) operator of an electron on site ${\mathbf{r}}=(x,y)$ of layer $z$ of spin $\sigma=\{\uparrow,\downarrow\}$. $\hat{n}_{z,{\mathbf{r}},\sigma}=\hat{c}_{z,{\mathbf{r}},\sigma}^{\dagger}\hat{c}_{z,{\mathbf{r}},\sigma}^{\phantom{\dagger}}$ is the corresponding particle number operator. Here the brakets $\langle z,z^{\prime}\rangle$ indicate neighboring layers along the $z$-axis while $\langle{\mathbf{r}}^{\phantom{\dagger}},{\mathbf{r}}^{\prime}\rangle$ refers to neighboring sites belonging to the same layer.
The central correlated region is under a time-periodic, homogeneous and monochromatic, electric field with frequency $\Omega$, which enters in the Hamiltonian 1 via the Peierls substitution [44] in the intralayer hopping as
$t_{z}(t)=t_{z}\ e^{-\mathrm{i}\frac{q}{\hbar}\left(\bm{r}-\bm{r}^{\prime}\right)\cdot\bm{A}(t)},$ (2)
where in Eq. 2 $\bm{A}$(t) is the time-dependent vector potential, $\hbar$ Planck’s constant and $q$ the charge of the electrons. We choose $\bm{A}(t)=A(t)\bm{e}_{0}$ to point along the lattice body diagonal of a hyper-cubic lattice $\bm{e}_{0}=(1,1,\dots,1)$, where $A(t)=\frac{\hbar}{qa}\mathscr{A}\sin(\Omega t)$ and $\mathscr{A}=-\frac{qE_{0}a}{\hbar\Omega}$, with $E_{0}$ is the electric field amplitude, and $a$ the lattice spacing [45, 46]. The electric field is then given in the temporal gauge by $\bm{E}=-\partial_{t}\bm{A}(t)=E_{0}\cos(\Omega t)\bm{e}_{0}$.
The second term in Eq. 1 describes the hopping of electrons in between layers with nearest-neighbor amplitude $t_{zz^{\prime}}$. The third term includes in the model the local onsite Hubbard interaction $U_{z}$ and the fourth one the onsite energies $\varepsilon_{z}^{(0)}$, the expression of which will be given in Sec. 3.1. $\hat{H}_{\mathrm{leads}}$ describes the Hamiltonian of the metallic leads, the details of which will be also given in Sec. 3.1.
@@
$\hat{H}_{\mathrm{e-ph}}=g\sum_{z,{\mathbf{r}},\sigma}\hat{n}_{z,{\mathbf{r}},\sigma}\hat{x}_{z,{\mathbf{r}}},$ (3)
where $g$ is the e-ph coupling constant and $\hat{x}_{z,{\mathbf{r}}}\equiv\frac{1}{\sqrt{2}}\left(\hat{b}^{\dagger}_{z,{\mathbf{r}}}+\hat{b}_{z,{\mathbf{r}}}\right)$, with $\hat{b}^{\dagger}_{z,{\mathbf{r}}}$ ($\hat{b}_{z,{\mathbf{r}}}$) which creates (annihilates) a phonon of the acoustic branch on site ${\mathbf{r}}=(x,y)$ of layer $z$, whose dispersion relation is described by $\hat{H}_{\mathrm{ph}}$, discussed in Sec 3.1.
@@
The xy-plane is modeled as a $d$-dimensional lattice in the limit $d\rightarrow\infty$ with the rescaling of the intralayer hopping as $t_{\parallel}=t_{\parallel}^{\ast}/(2\sqrt{d})$ [50]. Every momentum-dependent function depends on the electronic crystal momentum ${{\mathbf{k}}}$ only via $\epsilon=-2t_{\parallel}\sum_{i=1}^{d}\cos(k_{i}a)$ and $\overline{\epsilon}=-2t_{\parallel}\sum_{i=1}^{d}\sin(k_{i}a)$. Summations over the Brillouin zone are then performed using the joint density of states (JDOS) [45, 50] $\rho(\epsilon,\overline{\epsilon})=(1/\pi t_{\parallel}^{\ast 2})\exp[-(\epsilon^{2}+\overline{\epsilon}^{2})/t_{\parallel}^{\ast 2}]$.
In this paper we set $\hbar=k_{\mathrm{B}}=a=1=q$ and $t_{\parallel}^{\ast}/2=1$ as unit of energy.
3 Methods
3.1 Dyson equation
@@
$\underline{{\mathbf{G}}}^{-1}_{zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})$ $=\underline{{\mathbf{G}}}^{-1}_{0,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})-\underline{{\mathbf{\Sigma}}}_{zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})$ (4)
$-\underline{{\mathbf{\Sigma}}}_{\mathrm{e-ph},zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon}).$
The quantities in Eq. 4 are matrices in the Floquet representation, which we use to describe the periodic NESS of the system [25, 30, 31]. In the rest of the paper, we denote a Floquet-represented matrix as either $X_{mn}$ or $\mathbf{X}$. The underline in Eq. 4 stands for the Keldysh structure
$\underline{{\mathbf{X}}}\equiv\left(\begin{array}[]{cc}\mathbf{X}^{\mathrm{R}}&\mathbf{X}^{\mathrm{K}}\\ \mathbf{0}&\mathbf{X}^{\mathrm{A}}\end{array}\right)$ (5)
with $\textbf{X}^{\mathrm{R,A,K}}$ being the retarded, advanced and Keldysh components. We recall that $\textbf{X}^{\mathrm{A}}=(\textbf{X}^{\mathrm{R}})^{\dagger}$ and $\textbf{X}^{\mathrm{K}}\equiv\textbf{X}^{>}+\textbf{X}^{<}$, where $\textbf{X}^{\lessgtr}$ are the lesser and greater components [51, 52, 53, 54].
The GF which corresponds to the noninteracting terms in the Hamiltonian 1 reads
$\underline{G}^{-1}_{0,mn,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})=\underline{g}^{-1}_{0,mn,zz^{\prime}}(\omega_{n};\epsilon,\overline{\epsilon})$
$-\left[v^{2}_{\mathrm{l}}\underline{g}_{\mathrm{l}}(\omega_{n};\epsilon)\delta_{z,1}+v^{2}_{\mathrm{r}}\underline{g}_{\mathrm{r}}(\omega_{n};\epsilon)\delta_{z,\mathrm{L}}\right]\delta_{mn}\delta_{zz^{\prime}},$ (6)
@@
$\left[g^{-1}_{0}(\omega_{n};\epsilon,\overline{\epsilon}))\right]^{\mathrm{R}}_{mn,zz^{\prime}}=\left(\omega_{n}+\mathrm{i}0^{+}-\varepsilon_{z}\right)\delta_{mn}\delta_{zz^{\prime}}$
$-t_{zz^{\prime}}\delta_{mn}-\varepsilon_{mn}(\epsilon,\overline{\epsilon})\delta_{zz^{\prime}},$ (7)
@@
$\varepsilon_{z}=\varepsilon_{z}^{(0)}+\frac{\Phi}{2}-\frac{(z-1)\Phi}{(\mathrm{L}-1)},$ (8)
with $\varepsilon_{z}^{(0)}=-U/2$ at half-filling. The off-diagonal Floquet terms in Eq. 3.1 are given by the Floquet dispersion relation $\varepsilon_{mn}$ for a periodic field in a hyper-cubic lattice [45]
$\varepsilon_{mn}(\epsilon,\overline{\epsilon})=\left\{\begin{array}[]{ll}J_{m-n}(\mathscr{A})\ \epsilon&\quad m-n:\mathrm{even}\\ \mathrm{i}J_{m-n}(\mathscr{A})\ \overline{\epsilon}&\quad m-n:\mathrm{odd},\end{array}\right.$ (9)
where $J_{n}$ is the $n$-th order Bessel function of the first kind with argument $\mathscr{A}$ defined as in Sec. 2.
@@
$g^{\mathrm{R}}_{\rho}(\omega;\epsilon)$ $=\left(\omega-\varepsilon_{\rho}(\epsilon)+\mathrm{i}\gamma_{\rho}\right)^{-1},$
$g^{\mathrm{K}}_{\rho}(\omega;\epsilon)$ $=[g_{\rho}^{\mathrm{R}}(\omega;\epsilon)-g_{\rho}^{\mathrm{A}}(\omega;\epsilon)][1-2f(\omega,\mu_{\rho},\beta)],$ (10)
@@
The electron self-energy (SE) $\underline{\mathbf{\Sigma}}_{zz^{\prime}}$ and the e-ph SE $\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},zz^{\prime}}$ are obtained from DMFT [55, 56, 57] and its nonequibrium extension for driven inhomogeneous systems, i.e. real-space Floquet DMFT (F-DMFT) [45, 58, 33, 22, 36, 41, 42]. In this approximation, they are independent of the crystal momentum and spatially local, i.e. $\underline{{\mathbf{\Sigma}}}_{zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{\mathbf{\Sigma}}_{z}(\omega)\delta_{zz^{\prime}}$, $\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},z}(\omega)\delta_{zz^{\prime}}$. Details regarding real-space F-DMFT are given in A.
@@
$\Sigma_{\mathrm{e-ph},z}(z,z^{\prime})=\mathrm{i}g^{2}G_{\mathrm{loc},zz}(z,z^{\prime})D_{\mathrm{ph}}(z,z^{\prime}).$ (11)
One can easily extract the retarded and Keldysh components of the e-ph SE from Eq. 11, which are listed in Ref. [47].
In the current setup, away from half-filling, the Retarded component of $\underline{\underline{\underline{\Sigma}}}_{\mathrm{e-ph},z}(\omega)$ has also the contribution of the so-called Hartree shift $\varepsilon_{\mathrm{H},z}$, which enters in the $(0,0)$-Floquet matrix element as
$\Sigma_{\mathrm{e-ph},z,00}^{\mathrm{R}}(\omega)=\varepsilon_{\mathrm{H},z}+\Sigma_{\mathrm{e-ph},z,00,\mathrm{XC}}^{\mathrm{R}}(\omega),$ (12)
where $\Sigma_{\mathrm{e-ph},z,00,\mathrm{XC}}^{\mathrm{R}}(\omega)$ is the contribution coming from Eq. 11. The expression of $\varepsilon_{\mathrm{H},z}$ and its derivation are given in Sec. 4.1 and B respectively.
The Retarded component of the non-interacting phonon GF $\underline{D}_{\mathrm{ph}}(t,t^{\prime})$ is given by [61]
$D^{\mathrm{R}}_{\mathrm{ph}}(t,t^{\prime})=-\mathrm{i}\theta(t-t^{\prime})\int\mathop{}\!\mathrm{d}\omega\ \mathrm{e}^{-\mathrm{i}\omega\left(t-t^{\prime}\right)}A_{\mathrm{ph}}(\omega),$ (13)
where $A_{\mathrm{ph}}(\omega)=(\omega/\omega^{2}_{\mathrm{ph}})\mathrm{e}^{-|\omega|/\omega_{\mathrm{ph}}}$ is the acoustic phonons’ spectral function, with $\omega_{\mathrm{ph}}$ a soft cutoff frequency [62, 30]. The other Keldysh components are reconstructed via the fluctuation-dissipation theorem [63, 30].
3.2 Observables